In a capture-recapture study, a biologist tagged and released 320 deer. A month later she captured 40 deer, 20 of which had tags. About how many deer were in the population?
Answer by stanbon(75887)

In a capture-recapture study, a biologist tagged and released 320 deer. A month later she captured 40 deer, 20 of which had tags. About how many deer were in the population?
---
Use a proportion::
population/caught = # tagged/(tagged that are caught
----------------------
n/320 = 40/20
n = 320*2 = 640 (# in the population)
------------------
Cheers,
Stan H.
----------------